OpenStudy (anonymous):

A 100 W light bulb has an efficiency of 95%. How much thermal energy will this light bulb add to the inside of a room in 2.5 h if the bulb is on for the whole time?

OpenStudy (stormfire1):

\[1W=1\frac{J}{s}\]The 100W light bulb is 95% efficient so it adds 5W (or J/s) of thermal energy to the room. Converting 2.5 hours to seconds you get:\[2.5h*(\frac{3600s}{hr})=9000s\]And the final calculation:\[5\frac{J}{s}*9000s=45,000J\]